darren is 5 times as old as his sister emily, in ten years time the sum of their ages is 38 how old is emily
Answer by addingup(3677)

d = 5e
and
d+10+e+10 = 38 in this equation substitute for d per above equality:
5e+e+10+10 = 38; 6e+20 = 38; 6e = 18; e = 3 - Emily is 3 years old.
And Darren? (you may be asked)
d = 5e = 5(3) = 15 - Darren is 15